OpenStudy (anonymous):

Use Newton's method to find xsubscript6, when xsubscript0 = -7, for the function f(x) = x^2 + 13.5x + 45.

OpenStudy (anonymous):

Newton's method states \[x _{1}= x _{0} - f(x _{0})/f'(x _{0})\] So, to begin we shall evaluate the function for \[f(x _{0})\] = 45. We also need to find the first derivative and evaluate it at x = 0. \[f'(x) = 2x + 13.5\] So \[f'(x _{0})\] = 13.5 Substitution all that into Newton's method gives us x1 = 7 - 45/23.5 So x1 = 3 2/3 Now to continue we can use the same process over with the variation of the original formula\[x _{n+1} = x _{n} - f (x_{n})/f'(x _{n})\] It's a bit of work to get up to x6.